# Special pure gradings on simple Lie algebras of types $E_{6}$, $E_{7}$, $E_{8}$

### v1

@@ Abstract.
C.D. has been supported by the Spanish Ministerio de Ciencia e Innovación through grant PID2023-152673GB-I00 with FEDER funds; and by Junta de Andalucía through grant FQM-336.
@@ 1. Introduction
The purpose of this paper is to initiate the classification of gradings up to isomorphism on the simple Lie algebras of the $E$ types by classifying, both up to isomorphism and up to equivalence, the special pure gradings on these algebras. A group grading on a semisimple Lie algebra $\mathcal{L}$ over an algebraically closed field of characteristic zero is pure if at least one of its components, other than the identity component, contains a Cartan subalgebra of $\mathcal{L}$. This concept was introduced in [Hes82], motivated by the so-called Jordan gradings (see [Ale74, Eld09]). The particular symmetry of these gradings makes them especially useful for constructing manageable models of exceptional Lie algebras (for instance, the one in [Dr24]). A grading on $\mathcal{L}$ is special if its identity component is zero; these are precisely the almost fine gradings on $\mathcal{L}$ that have toral rank 00.
@@ 2.1. Quadratic forms in characteristic 2
In characteristic 2, a hyperbolic plane $H$ is spanned by $u,v\in V$ with $B_{Q}(u,v)=1$ and $Q(u)=Q(v)=0$. The maximum $r$ such that $V=H_{1}\obot H_{2}\obot\dots\obot H_{r}\obot W$, with each $H_{i}$ a hyperbolic plane, is called the Witt index, and coincides with the dimension of a maximal totally isotropic subspace of $V$ (in fact, any two of these are isometric, due to Witt’s Extension Theorem). The Witt index equals $m$ in case (1), and $m$ and $m-1$, respectively, in cases (2a) and (2b).
@@ 2.2. Weyl groups of Lie algebras of the $E$ types
We will now recall why the Weyl groups of the simple Lie algebras of types $E_{6}$, $E_{7}$ and $E_{8}$ are closely related to orthogonal groups over the field $\mathbb{F}_{2}$. Let $\mathcal{L}$ be one of these Lie algebras over an algebraically closed field $\mathbb{F}$ of characteristic 00, let $\mathcal{H}$ be a Cartan subalgebra of $\mathcal{L}$, and $\Phi$ the root system associated to $\mathcal{H}$. That is, we have the root space decomposition
@@
where $\mathcal{L}_{\alpha}=\{x\in\mathcal{L}\mid[h,x]=\alpha(h)x\;\forall h\in\mathcal{H}\}$. Let (.|.)(.\,|\,.)( . | . ) be the bilinear form on $\mathcal{H}^{*}$, dual to the bilinear form obtained from the restriction of the Killing form to $\mathcal{H}$, normalized so that $(\alpha\,|\,\alpha)=2$ for any $\alpha\in\Phi$. This bilinear form restricts to a $\mathbb{Z}$-bilinear form on the root lattice $R=\mathbb{Z}\Phi$. Consider the associated $\mathbb{Z}$-valued quadratic form $\frac{1}{2}(r\,|\,r)$, whose polarization is (.|.)(.\,|\,.)( . | . ). This quadratic form induces a quadratic form on the vector space $\overline{R}:=R/2R$ over the field of two elements $\mathbb{F}_{2}$:
@@
with polarization $b_{q}$ induced by the bilinear form (.|.)(.\,|\,.)( . | . ). Note that $q(\bar{\alpha})=1$ for any root $\alpha\in\Phi$. Also, in the cases at hand, if $\beta\neq\pm\alpha$ then $\bar{\beta}\neq\bar{\alpha}$.
@@
Let $\mathcal{L}$ be a simple Lie algebra of type $E_{6}$, $E_{7}$, or $E_{8}$ over an algebraically closed field $\mathbb{F}$ of characteristic 00. Then the image of the homomorphism $\rho$ is the orthogonal group $\mathrm{O}(q):=\mathrm{O}(\bar{R},q)$, and the kernel is {±$\{\pm I\}$. Thus, the Weyl group is isomorphic to
@@
Any automorphism of the root system $\Phi$ preserves the quadratic form $\frac{1}{2}(.\,|\,.)$ on the root lattice $R=\mathbb{Z}\Phi$, and hence the image of $\rho$ is contained in $\mathrm{O}(q)$. Now, the Weyl group is generated by the reflections relative to roots, and $\rho$ maps them to the reflections (sometimes called orthogonal transvections in characteristic $2$) relative to the nonisotropic vectors of $\overline{R}$. Recall that Cartan-Dieudonné Theorem [Die48] says that, with one exception, reflections generate the orthogonal group of a regular quadratic form. To be precise, for even-dimensional regular quadratic forms over a field of characteristic $2$, the orthogonal group is generated by reflections unless the dimension is $4$, the Arf invariant is 00 and the ground field is $\mathbb{F}_{2}$ (see, e.g., [Gro02, Theorem 14.16]), and we have dimensions $6$ and $8$ in our cases. In odd dimension, the orthogonal group is isomorphic to a symplectic group, so the orthogonal group can be shown to be generated by reflections using the fact that the symplectic group is generated by transvections ([Gro02, Theorem 3.4]). This completes the proof of the surjectivity of $\rho$.
If $\mu\in\operatorname{\mathrm{Aut}}(\Phi)$ induces the identity map on $\overline{R}$, then it must map each root $\alpha$ to ±$\pm\alpha$, and it is clear that the sign must be the same for all simple roots, because the Dynkin diagram is connected. Hence $\ker\rho=\{\pm I\}$.
Finally, the Weyl group is $\operatorname{\mathrm{Aut}}(\Phi)$ for types $E_{7}$ and $E_{8}$, whereas for type $E_{6}$ the Weyl group is a complement of {±$\{\pm I\}$ in $\operatorname{\mathrm{Aut}}(\Phi)$. Also, for type $E_{7}$, $\operatorname{\mathrm{Aut}}(\Phi)$ is the direct product of {±$\{\pm I\}$ and the subgroup of elements of determinant $1$, as transformations of the euclidean vector space spanned by the roots. ∎
@@ 2.3. Gradings on Lie algebras
For any $(s_{1},s_{2})\in S_{*}$, denote by $s_{1}*s_{2}$ the unique element of $S$ such that $[\mathcal{L}_{s_{1}},\mathcal{L}_{s_{2}}]\subset\mathcal{L}_{s_{1}*s_{2}}$.
@@
$\Gamma^{*}:\mathcal{L}=\bigoplus_{(g,\lambda)\in G\times\widehat{\mathcal{T}}}\mathcal{L}_{(g,\lambda)},\textrm{ for }\mathcal{L}_{(g,\lambda)}:=\{x\in\mathcal{L}_{g}\mid\tau(x)=\lambda(\tau)x\ \forall\tau\in\mathcal{T}\}.$ (2.5)
@@
Let $\mathcal{L}$ be a semisimple Lie algebra of rank $l$. Consider a Cartan subalgebra $\mathcal{H}$ of $\mathcal{L}$, the root system $\Phi$ associated to $\mathcal{H}$, and the root space decomposition as in (2.1). As in Example 2.9, there exists an order $2$ automorphism $\sigma$ of $\mathcal{L}$ which is $-\mathrm{id}$ on $\mathcal{H}$, and hence takes any root space $\mathcal{L}_{\alpha}$ to $\mathcal{L}_{-\alpha}$. Then $\sigma$ commutes with the elements of order $2$ in the torus $\mathcal{T}$ associated to $\mathcal{H}$. The eigenspaces of the quasitorus $\mathcal{T}_{2}\times\langle\sigma\rangle$, where $\mathcal{T}_{2}:=\{\tau\in\mathcal{T}\mid\tau^{2}=\mathrm{id}\}$, are the homogeneous components of a special pure grading on $\mathcal{L}$, whose universal group is isomorphic to $\mathbb{Z}_{2}^{l+1}$.
@@ 3. Weyl groups of special pure gradings
This section will be devoted to describe the special pure gradings of semisimple Lie algebras and to begin our study of the Weyl groups, following the work of Hesselink [Hes82].
Let $\mathcal{L}$ be a finite-dimensional semisimple Lie algebra over an algebraically closed ground field $\mathbb{F}$ of characteristic 00, and let
@@
For the sake of completeness, we include a sketch of the proof of (3.1). If $\chi$ is a character of $G$ with $\chi(s)\neq 1$, the automorphism $\sigma=\tau_{\chi}$ restricts to $\chi(s)\mathrm{id}$ on $\mathcal{L}_{s}$ and hence on $\mathcal{H}$. This forces $\chi(s)^{-1}\alpha$ to be a root for any root $\alpha$ relative to $\mathcal{H}$. It follows that $\chi(s)=-1$, which forces $\sigma^{2}=\mathrm{id}$ ([Eld22, Lemma 2.1]). This same argument shows that $\chi(s)=1$ or $-1$ for any $\chi\in\widehat{G}$, and hence $\operatorname{\mathrm{Diag}}(\Gamma)$ is contained in the centralizer of $\sigma$ in $\{\varphi\in\operatorname{\mathrm{Aut}}(\mathcal{L})\mid\varphi|_{\mathcal{H}}=\pm\mathrm{id}\}=\mathcal{T}\rtimes\langle\sigma\rangle$, and this centralizer is $\mathcal{T}_{2}\times\langle\sigma\rangle$ ([Eld22, Lemma 2.1]). It follows that $\operatorname{\mathrm{Diag}}(\Gamma)=\bigl{(}\operatorname{\mathrm{Diag}}(\Gamma)\cap\mathcal{T}_{2}\bigr{)}\times\langle\sigma\rangle$.
@@
Recall that, for $\mathcal{L}$ of types $E_{6}$, $E_{7}$, $E_{8}$, we have defined in (2.2) a regular quadratic form $q$ on $\overline{R}$, regarded as a vector space over $\mathbb{F}_{2}$. Since $\mathcal{T}_{2}$ is identified with the character group of $\overline{R}$, which can be considered as the dual space of $\overline{R}$, we can transport the quadratic form $q$ in the cases $E_{6}$ and $E_{8}$ along the isomorphism $\overline{R}\to\mathcal{T}_{2}$ defined by the nondegenerate bilinear form $b_{q}$. The resulting quadratic form $\hat{q}$ on $\mathcal{T}_{2}$ has the following meaning. For $\mathcal{L}$ of type $E_{8}$, there are two conjugacy classes of automorphisms of order $2$, which are distinguished by the subalgebra of fixed points, and $\hat{q}$ takes value 00 on the nontrivial elements $\tau\in\mathcal{T}_{2}$ for which $\mathcal{L}^{\tau}$ is of type $D_{8}$ and value $1$ on those for which $\mathcal{L}^{\tau}$ is of type $E_{7}+A_{1}$. Moreover, the elements of $\mathcal{T}_{2}\times\langle\sigma\rangle$ that are not in $\mathcal{T}_{2}$ give type $D_{8}$. For $\mathcal{L}$ of type $E_{6}$, there are two conjugacy classes of inner automorphisms of order $2$, and $\hat{q}$ takes value 00 when $\mathcal{L}^{\tau}$ is of type $D_{5}+Z$ (where $Z$ stands for the $1$-dimensional center) and value $1$ when $\mathcal{L}^{\tau}$ is of type $A_{5}+A_{1}$. Moreover, the elements of $\mathcal{T}_{2}\times\langle\sigma\rangle$ that are not in $\mathcal{T}_{2}$ are outer automorphisms (with $\mathcal{L}^{\tau}$ of type $C_{4}$).
@@
where, for any $g\in G$, $\bar{g}$ denotes its coset $g\langle s\rangle$, and $\mathcal{L}_{\bar{g}}=\mathcal{L}_{g}\oplus\mathcal{L}_{gs}$. As the Cartan subalgebra $\mathcal{H}$ is contained in the neutral homogeneous component $\mathcal{L}_{\bar{s}}=\mathcal{L}_{\bar{e}}$, the grading $\overline{\Gamma}$ is a coarsening of the Cartan grading (2.1) and this gives a surjective group homomorphism
@@
¯$\begin{split}\overline{\mathcal{L}}_{r+E}&=\bigoplus_{\alpha\in\Phi^{+}\cap(r+E)}\left(\mathcal{L}_{\alpha}\oplus\mathcal{L}_{-\alpha}\right),\qquad\text{if$r\not\in E$,}\\ \overline{\mathcal{L}}_{E}&=\mathcal{H}\oplus\left(\bigoplus_{\alpha\in\Phi^{+}\cap E}\left(\mathcal{L}_{\alpha}\oplus\mathcal{L}_{-\alpha}\right)\right).\end{split}$ (3.7)
@@
Let $\Gamma$ be a pure grading on a semisimple Lie algebra $\mathcal{L}$ over an algebraically closed field of characteristic 00, with universal group $G$. Let $s\in G$ be an element such that $\mathcal{L}_{s}$ contains a Cartan subalgebra $\mathcal{H}$. Let $\Phi$ be the root system relative to $\mathcal{H}$, $R=\mathbb{Z}\Phi$ the root lattice, and $E$ the complementary lattice of $\Gamma$. Then $\Gamma$ is special if and only if $\Phi\cap E=\emptyset$.
@@
In many situations, there is a unique $s\in G$ containing a Cartan subalgebra, so $W_{s}(\Gamma)$ is the whole Weyl group $W(\Gamma)$.
@@
$\left(\begin{array}[]{c|c}\quad I_{l}&\begin{matrix}0\\[-3.0pt] \vdots\\ 0\end{matrix}\\ \hline\cr\\[-8.0pt] \xi&1\end{array}\right),$
@@
where the first map is the natural homomorphism and the second is the inverse of the isomorphism $\pi_{E}$ in (3.3). It induces a surjective group homomorphim
@@
Let $\Gamma$ be a special pure grading on the semisimple Lie algebra $\mathcal{L}$, over an algebraically closed field of characteristic 00, with universal group $G$. Let $s\in G$ be an element such that $\mathcal{L}_{s}$ contains a Cartan subalgebra $\mathcal{H}$. Then, with the notations above, the stabilizer of $s$ in the Weyl group $W(\Gamma)$ is obtained as follows:
@@
$\left(\begin{array}[]{c|c}\quad\mu_{\varphi}^{\prime}&\begin{matrix}0\\[-3.0pt] \vdots\\ 0\end{matrix}\\ \hline\cr\\[-8.0pt] \xi&1\end{array}\right),$
@@
$\left(\begin{array}[]{c|c}\quad\mu^{\prime}&\begin{matrix}0\\[-3.0pt] \vdots\\ 0\end{matrix}\\ \hline\cr\\[-8.0pt] \xi&1\end{array}\right),$

### v2

@@ Abstract.
C.D. has been supported by the Spanish Ministerio de Ciencia e Innovación through project PID2023-152673NB-I00 and by the Junta de Andalucía through project PPRO-FQM336-G-2023, all of them with FEDER funds
@@ 1. Introduction
The purpose of this paper is to initiate the classification of gradings up to isomorphism on the simple Lie algebras of the $E$ types by classifying, both up to isomorphism and up to equivalence, the special pure gradings on these algebras. A group grading on a semisimple Lie algebra $\mathcal{L}$ over an algebraically closed field of characteristic zero is pure if at least one of its components, other than the identity component, contains a Cartan subalgebra of $\mathcal{L}$. This concept was introduced in [Hes82], motivated by the so-called Jordan gradings (see [Ale74, Eld09]). The particular symmetry of these gradings makes them especially useful for constructing manageable models of exceptional Lie algebras (for instance, the one in [Dr24]). A grading on $\mathcal{L}$ is special if its identity component is zero; these are precisely the almost fine gradings on $\mathcal{L}$ that have toral rank 0.
@@ 2.1. Quadratic forms in characteristic 2
In characteristic 2, a hyperbolic plane $H$ is spanned by $u,v\in V$ with $B_{Q}(u,v)=1$ and $Q(u)=Q(v)=0$. The maximum $r$ such that $V=H_{1}\obot H_{2}\obot\dots\obot H_{r}\obot W$, with each $H_{i}$ a hyperbolic plane, is called the Witt index, and coincides with the dimension of a maximal totally isotropic subspace of $V$ (in fact, any two of these are isometric, due to Witt’s Extension Theorem). The Witt index equals $m$ in case (1), and $m$ and m1m-1, respectively, in cases (2a) and (2b).
@@ 2.2. Weyl groups of Lie algebras of the $E$ types
We will now recall why the Weyl groups of the simple Lie algebras of types $E_{6}$, $E_{7}$ and $E_{8}$ are closely related to orthogonal groups over the field $\mathbb{F}_{2}$. Let $\mathcal{L}$ be one of these Lie algebras over an algebraically closed field $\mathbb{F}$ of characteristic 0, let $\mathcal{H}$ be a Cartan subalgebra of $\mathcal{L}$, and $\Phi$ the root system associated to $\mathcal{H}$. That is, we have the root space decomposition
@@
where $\mathcal{L}_{\alpha}=\{x\in\mathcal{L}\mid[h,x]=\alpha(h)x\;\forall h\in\mathcal{H}\}$. Let (.|.)(.\,|\,.) be the bilinear form on $\mathcal{H}^{*}$, dual to the bilinear form obtained from the restriction of the Killing form to $\mathcal{H}$, normalized so that $(\alpha\,|\,\alpha)=2$ for any $\alpha\in\Phi$. This bilinear form restricts to a $\mathbb{Z}$-bilinear form on the root lattice $R=\mathbb{Z}\Phi$. Consider the associated $\mathbb{Z}$-valued quadratic form $\frac{1}{2}(r\,|\,r)$, whose polarization is (.|.)(.\,|\,.). This quadratic form induces a quadratic form on the vector space $\overline{R}:=R/2R$ over the field of two elements $\mathbb{F}_{2}$:
@@
with polarization $b_{q}$ induced by the bilinear form (.|.)(.\,|\,.). Note that $q(\bar{\alpha})=1$ for any root $\alpha\in\Phi$. Also, in the cases at hand, if $\beta\neq\pm\alpha$ then $\bar{\beta}\neq\bar{\alpha}$.
@@
Let $\mathcal{L}$ be a simple Lie algebra of type $E_{6}$, $E_{7}$, or $E_{8}$ over an algebraically closed field $\mathbb{F}$ of characteristic 0. Then the image of the homomorphism $\rho$ is the orthogonal group $\mathrm{O}(q):=\mathrm{O}(\bar{R},q)$, and the kernel is $\{\pm I\}$. Thus, the Weyl group is isomorphic to
@@
Any automorphism of the root system $\Phi$ preserves the quadratic form $\frac{1}{2}(.\,|\,.)$ on the root lattice $R=\mathbb{Z}\Phi$, and hence the image of $\rho$ is contained in $\mathrm{O}(q)$. Now, the Weyl group is generated by the reflections relative to roots, and $\rho$ maps them to the reflections (sometimes called orthogonal transvections in characteristic $2$) relative to the nonisotropic vectors of $\overline{R}$. Recall that Cartan-Dieudonné Theorem [Die48] says that, with one exception, reflections generate the orthogonal group of a regular quadratic form. To be precise, for even-dimensional regular quadratic forms over a field of characteristic $2$, the orthogonal group is generated by reflections unless the dimension is $4$, the Arf invariant is 0 and the ground field is $\mathbb{F}_{2}$ (see, e.g., [Gro02, Theorem 14.16]), and we have dimensions $6$ and $8$ in our cases. In odd dimension, the orthogonal group is isomorphic to a symplectic group, so the orthogonal group can be shown to be generated by reflections using the fact that the symplectic group is generated by transvections ([Gro02, Theorem 3.4]). This completes the proof of the surjectivity of $\rho$.
If $\mu\in\operatorname{\mathrm{Aut}}(\Phi)$ induces the identity map on $\overline{R}$, then it must map each root $\alpha$ to $\pm\alpha$, and it is clear that the sign must be the same for all simple roots, because the Dynkin diagram is connected. Hence $\ker\rho=\{\pm I\}$.
Finally, the Weyl group is $\operatorname{\mathrm{Aut}}(\Phi)$ for types $E_{7}$ and $E_{8}$, whereas for type $E_{6}$ the Weyl group is a complement of $\{\pm I\}$ in $\operatorname{\mathrm{Aut}}(\Phi)$. Also, for type $E_{7}$, $\operatorname{\mathrm{Aut}}(\Phi)$ is the direct product of $\{\pm I\}$ and the subgroup of elements of determinant $1$, as transformations of the euclidean vector space spanned by the roots. ∎
@@ 2.3. Gradings on Lie algebras
For any $(s_{1},s_{2})\in S_{*}$, denote by $s_{1}*s_{2}$ the unique element of $S$ such that $\mathcal{L}_{s_{1}},\mathcal{L}_{s_{2}}]\subset\mathcal{L}_{s_{1}*s_{2}}$.
@@
$\Gamma^{*}:\mathcal{L}=\hskip-12.0pt\bigoplus_{(g,\lambda)\in G\times\widehat{\mathcal{T}}}\hskip-12.0pt\mathcal{L}_{(g,\lambda)},\textrm{ for }\mathcal{L}_{(g,\lambda)}:=\{x\in\mathcal{L}_{g}\mid\tau(x)=\lambda(\tau)x\ \forall\tau\in\mathcal{T}\}.$ (2.5)
@@
Let $\mathcal{L}$ be a semisimple Lie algebra of rank $l$. Consider a Cartan subalgebra $\mathcal{H}$ of $\mathcal{L}$, the root system $\Phi$ associated to $\mathcal{H}$, and the root space decomposition as in (2.1). As in Example 2.9, there exists an order $2$ automorphism $\sigma$ of $\mathcal{L}$ which is $\mathrm{id}$ on $\mathcal{H}$, and hence takes any root space $\mathcal{L}_{\alpha}$ to $\mathcal{L}_{-\alpha}$. Then $\sigma$ commutes with the elements of order $2$ in the torus $\mathcal{T}$ associated to $\mathcal{H}$. The eigenspaces of the quasitorus $\mathcal{T}_{2}\times\langle\sigma\rangle$, where $\mathcal{T}_{2}:=\{\tau\in\mathcal{T}\mid\tau^{2}=\mathrm{id}\}$, are the homogeneous components of a special pure grading on $\mathcal{L}$, whose universal group is isomorphic to $\mathbb{Z}_{2}^{l+1}$.
@@ 3. Weyl groups of special pure gradings
This section will be devoted to describing the special pure gradings of semisimple Lie algebras and to begin our study of the Weyl groups, following the work of Hesselink [Hes82].
Let $\mathcal{L}$ be a finite-dimensional semisimple Lie algebra over an algebraically closed ground field $\mathbb{F}$ of characteristic 0, and let
@@
For the sake of completeness, we include a sketch of the proof of (3.1). If $\chi$ is a character of $G$ with $\chi(s)\neq 1$, the automorphism $\sigma=\tau_{\chi}$ restricts to $\chi(s)\mathrm{id}$ on $\mathcal{L}_{s}$ and hence on $\mathcal{H}$. This forces $\chi(s)^{-1}\alpha$ to be a root for any root $\alpha$ relative to $\mathcal{H}$. It follows that $\chi(s)=-1$, which forces $\sigma^{2}=\mathrm{id}$ ([Eld22, Lemma 2.1]). This same argument shows that $\chi(s)=1$ or 1-1 for any $\chi\in\widehat{G}$, and hence $\operatorname{\mathrm{Diag}}(\Gamma)$ is contained in the centralizer of $\sigma$ in $\{\varphi\in\operatorname{\mathrm{Aut}}(\mathcal{L})\mid\varphi|_{\mathcal{H}}=\pm\mathrm{id}\}=\mathcal{T}\rtimes\langle\sigma\rangle$, and this centralizer is $\mathcal{T}_{2}\times\langle\sigma\rangle$ ([Eld22, Lemma 2.1]). It follows that $\operatorname{\mathrm{Diag}}(\Gamma)=\bigl(\operatorname{\mathrm{Diag}}(\Gamma)\cap\mathcal{T}_{2}\bigr)\times\langle\sigma\rangle$.
@@
Recall that, for $\mathcal{L}$ of types $E_{6}$, $E_{7}$, $E_{8}$, we have defined in (2.2) a regular quadratic form $q$ on $\overline{R}$, regarded as a vector space over $\mathbb{F}_{2}$. Since $\mathcal{T}_{2}$ is identified with the character group of $\overline{R}$, which can be considered as the dual space of $\overline{R}$, we can transport the quadratic form $q$ in the cases $E_{6}$ and $E_{8}$ along the isomorphism $\overline{R}\to\mathcal{T}_{2}$ defined by the nondegenerate bilinear form $b_{q}$. The resulting quadratic form $\hat{q}$ on $\mathcal{T}_{2}$ has the following meaning. For $\mathcal{L}$ of type $E_{8}$, there are two conjugacy classes of automorphisms of order $2$, which are distinguished by the subalgebra of fixed points, and $\hat{q}$ takes value 0 on the nontrivial elements $\tau\in\mathcal{T}_{2}$ for which $\mathcal{L}^{\tau}$ is of type $D_{8}$ and value $1$ on those for which $\mathcal{L}^{\tau}$ is of type $E_{7}+A_{1}$. Moreover, the elements of $\mathcal{T}_{2}\times\langle\sigma\rangle$ that are not in $\mathcal{T}_{2}$ give type $D_{8}$. For $\mathcal{L}$ of type $E_{6}$, there are two conjugacy classes of inner automorphisms of order $2$, and $\hat{q}$ takes value 0 when $\mathcal{L}^{\tau}$ is of type $D_{5}+Z$ (where $Z$ stands for the $1$-dimensional center) and value $1$ when $\mathcal{L}^{\tau}$ is of type $A_{5}+A_{1}$. Moreover, the elements of $\mathcal{T}_{2}\times\langle\sigma\rangle$ that are not in $\mathcal{T}_{2}$ are outer automorphisms (with $\mathcal{L}^{\tau}$ of type $C_{4}$).
@@
where, for any $\in G, g¯\bar{g}$ denotes its coset $g\langle s\rangle$, and $\mathcal{L}_{\bar{g}}=\mathcal{L}_{g}\oplus\mathcal{L}_{gs}$. As the Cartan subalgebra $\mathcal{H}$ is contained in the neutral homogeneous component $\mathcal{L}_{\bar{s}}=\mathcal{L}_{\bar{e}}$, the grading $\overline{\Gamma}$ is a coarsening of the Cartan grading (2.1) and this gives a surjective group homomorphism
@@
$\begin{split}\overline{\mathcal{L}}_{r+E}&=\bigoplus_{\alpha\in\Phi^{+}\cap(r+E)}\left(\mathcal{L}_{\alpha}\oplus\mathcal{L}_{-\alpha}\right),\qquad\text{if $r\not\in E$,}\\ \overline{\mathcal{L}}_{E}&=\mathcal{H}\oplus\left(\bigoplus_{\alpha\in\Phi^{+}\cap E}\left(\mathcal{L}_{\alpha}\oplus\mathcal{L}_{-\alpha}\right)\right).\end{split}$ (3.7)
@@
Let $\Gamma$ be a pure grading on a semisimple Lie algebra $\mathcal{L}$ over an algebraically closed field of characteristic 0, with universal group $G$. Let $s\in G$ be an element such that $\mathcal{L}_{s}$ contains a Cartan subalgebra $\mathcal{H}$. Let $\Phi$ be the root system relative to $\mathcal{H}$, $R=\mathbb{Z}\Phi$ the root lattice, and $E$ the complementary lattice of $\Gamma$. Then $\Gamma$ is special if and only if $\Phi\cap E=\emptyset$.
@@
In many situations, there is a unique $s\in G$ with $L_{s}$ containing a Cartan subalgebra, so $W_{s}(\Gamma)$ is the whole Weyl group $W(\Gamma)$.
@@
$\left(\begin{array}[]{c|c}\quad I_{l}\quad&\begin{matrix}0\\[-3.0pt] \vdots\\ 0\end{matrix}\\ \hline\cr\\[-8.0pt] \xi&1\end{array}\right),$
@@
where the first map is the natural homomorphism and the second is the inverse of the isomorphism $\pi_{E}$ in (3.3). It induces a surjective group homomorphism
@@
Let $\Gamma$ be a special pure grading on the semisimple Lie algebra $\mathcal{L}$, over an algebraically closed field of characteristic 0, with universal group $G$. Let $s\in G$ be an element such that $\mathcal{L}_{s}$ contains a Cartan subalgebra $\mathcal{H}$. Then, with the notations above, the stabilizer of $s$ in the Weyl group $W(\Gamma)$ is obtained as follows:
@@
$\left(\begin{array}[]{c|c}\quad\mu_{\varphi}^{\prime}\quad&\begin{matrix}0\\[-3.0pt] \vdots\\ 0\end{matrix}\\ \hline\cr\\[-8.0pt] \xi&1\end{array}\right),$
@@
$\left(\begin{array}[]{c|c}\quad\mu^{\prime}\quad&\begin{matrix}0\\[-3.0pt] \vdots\\ 0\end{matrix}\\ \hline\cr\\[-8.0pt] \xi&1\end{array}\right),$